solve the system by substitution:
4x-4y=0
y=2x-5

Answer by checkley71(8403)

Y=2X-5
4X-4(2X-5)=0
4X-8X+20=0
-4X=-20
X=-20/-4
X=5 ANSWER.
Y=2*5-5
Y=10-5
Y=5
PROOF
4*5-4*5=0
20-20=0
0=0